# Quantifying the Dynamics of Harm Caused by Retracted Research

### v1

@@ Abstract
Despite enormous efforts devoted to understand the characteristics and impacts of retracted papers, little is known about the mechanisms underlying the dynamics of their harm and the dynamics of its propagation. Here, we propose a citation-based framework to quantify the harm caused by retracted papers, aiming to uncover why their harm persists and spreads so widely. We uncover an “attention escape” mechanism, wherein retracted papers postpone significant harm, more prominently affect indirectly citing papers, and inflict greater harm on citations in journals with an impact factor less than . This mechanism allows retracted papers to inflict harm outside the attention of authors and publishers, thereby evading their intervention. This study deepens understanding of the harm caused by retracted papers, emphasizes the need to activate and enhance the attention of authors and publishers, and offers new insights and a foundation for strategies to mitigate their harm and prevent its spread.
@@
As of December 26, 2024, nearly 60,000 papers have been publicly retracted, and this staggering number may only represent the tip of the iceberg (?, ?). These papers can even deal a fatal blow to entire research fields, as exemplified by the devastating cases of fraud in stem cell research (?) and Alzheimer’s research (?). Worse still, in recent years, the number of retracted papers has consistently been remarkably high each year, making the challenges we face increasingly severe (?, ?, ?).
@@
Research on retracted papers has made significant progress, but most efforts have focused on analyzing the characteristics of retracted papers (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?), with relatively little attention given to their harm and mechanisms of dissemination (?, ?). Currently, researchers widely believe that the persistent harm of retracted papers stems from the difficulty authors, reviewers, and journals face in identifying and tracking their citations (?, ?, ?). To address this issue, tools such as Feet of Clay Detector (?), Annulled Detector (?), and RetractoBot (?) have been developed to flag citations of retracted papers and alert authors to handle such references with caution. However, recent studies suggest that these measures have been less effective than anticipated. One study found that the act of flagging citations of retracted papers provoked resistance from some authors (?). Another study flagged 88 articles citing retracted papers and sent notification emails to the authors, but only 44 responded, and just 11 implemented corrective actions (?). These findings indicate that current understanding of the harm caused by retracted papers remains significantly limited and biased. To mitigate the long-term impact of retracted literature, more in-depth research and fresh perspectives are urgently needed to drive meaningful progress in this field.
@@
Citations are a key and typical form through which the harm of retracted papers arises and spreads. At the same time, citations, like publications, are regarded as the “currency” of academia, forming the foundation of scholarly influence (?). They bolster the reputation of authors and institutions, drive career advancement, secure funding opportunities, and contribute to journal rankings and impact metrics (?, ?, ?, ?). Recognizing this dual role, we propose a citation-based quantitative framework to deepen understanding of the harm caused by retracted papers and uncover the mechanisms driving it (see “Materials and Methods” in the Supplementary Materials for details). To this end, we analyzed a refined dataset of 210,309,367 publications across 23 disciplines, extracted from the Semantic Scholar dataset following data cleaning and preprocessing.(, see “Materials and Methods” in the Supplementary Materials) (?). By integrating data from the Retraction Watch dataset () (?), which, after data cleaning and preprocessing, catalogs 39,449 retracted papers, we constructed an extensive citation network of retracted publications. This network enabled us to systematically trace the pathways through which harm originates and propagates, uncovering patterns that link retracted papers to their broader impact on the research community.
@@
As illustrated in Figure 1, to investigate the long-term impact of retracted papers, we focused on papers published before 2014 within the , creating a subset dataset named to track 10 years of citation dynamics. From the dataset, we identified papers directly citing those in the to form the citation distance 1 set (). Subsequently, we iteratively extracted papers citing to form the citation distance 2, 3, 4, 5, and 6 sets, denoted as , , , , and , respectively. For each of these sets, we calculated annual citation counts over 10 years and determined the expected citation counts based on contemporaneous papers published in the same journals and fields. The shortfall between observed and expected citations was defined as the harm caused by retracted papers, providing a quantitative measure of their harm and the dissemination of the harm. Details of the methodology are provided in the “Materials and Methods” section of the Supplementary Materials. Our research uncovers an “attention escape” mechanism by which retracted papers cause long-term, widespread, and significant harm to the scientific community. This mechanism operates through three key strategies: (1) Papers citing retracted studies experience insignificant harm during the early years when they receive the most attention, but this harm becomes increasingly significant over time. (2) Papers that indirectly cite retracted papers suffer more significant harm compared to those that directly cite them, and the severity of this harm increases with greater citation distance. (3) Papers citing retracted papers that are published in journals with an impact factor below 10 experience more significant harm compared to those published in higher-impact, more widely recognized journals. This mechanism provides a new understanding of the harm caused by retracted papers, prompting us to shift our focus toward reactivating and enhancing the attention of authors, journals, and other stakeholders (?), thereby offering a fresh perspective for exploring new measures to mitigate the harm of retracted papers.
Results
@@
Temporal Dynamics of Harm Arising from Retracted Papers.
@@
To investigate how the harm of retracted papers evolves over time, we analyzed the annual citation impact on papers citing these retracted works within 10 years after their publication. Notably, the ”10 years after publication” refers to the period starting from the second year through to the eleventh year after the citing papers were published. The first year was excluded from the analysis because publication dates can vary widely throughout the year (e.g., January or December), which may result in significant biases in citation data for the first year. As shown in Figure 2(A), the harm to papers directly citing retracted papers during the first four years after publication was not significant, with the highest median (interquartile range) harm observed in the fourth year, at 18.67% (-62.92% to 68.09%). Starting in the fifth year, the harm to papers directly citing retracted papers exceeded 20%, with a median (interquartile range) of 22.84% (-57.55% to 71.98%). The lowest median (interquartile range) damage to the paper directly citing retracted papers was recorded in the second year at 10.68% (-68.71% to 62.46%). Beyond the second year, the harm to the paper directly citing retracted papers increased annually, although the rate of increase slowed over time, reaching a median (interquartile range) of 40.08% (-41.39% to 100%) by the tenth year. These findings suggest that the harm caused by retracted papers to citing papers accumulates over time, making long-term impacts unavoidable. We further analyzed the field-specific effects of retracted papers using Semantic Scholar’s discipline classifications. As shown in Figure 2(B), disciplines with more than 300 citing articles (representing 99.65% of the dataset) exhibited the harm to the paper directly citing retracted papers trends similar to the overall pattern. However, certain fields, computer, physics, business, and economics (representing 2.26% of the dataset), showed significant harm to the paper directly citing retracted papers earlier in the citing papers’ lifespans.
It should be noted that a single article may belong to multiple fields, so the total percentages exceed 100%. Detailed statistics for each discipline are provided in Supplementary Table S1.
@@
Notably, in exploring the reasons behind the lack of attention to the harmful effects of problematic papers, we identified two phenomena potentially linked to the insufficient focus on the damaging mechanisms we uncovered, based on the interests most relevant to stakeholders. Two primary actors contribute to the dissemination of damage: authors (who cite retracted papers) and journals (which publish papers citing retracted works). In our investigation of authors, we focused on universities as key research institutions. As shown in Table Citation Timing Analysis of Harm., among the top 30 universities in the world, we found that 28 institutions place significant emphasis on recent research outputs (typically within the past 3 to 5 years) when considering faculty promotions (?). In our investigation of journals, we found that most journals prioritize impact factor as a key metric of interest, with impact factor calculations typically based on papers published within the past 3 years (?). These phenomena suggest that papers from earlier years receive far less attention from stakeholders, which could be a key reason why the potential harm of problematic papers is overlooked.
@@
Characteristics of the Propagation of Harm Arising from Retracted Papers.
To explore how the harm from retracted papers propagates, we drew upon the ”six degrees of separation” theory to study articles that are connected to retracted papers via indirect citations within six steps (?). As shown in Figure 3(A), while the pattern of harm experienced by indirectly citing articles is similar to that of directly citing articles, the harm to indirectly citing articles is more pronounced. Specifically, within the first four years after publication ( 2nd year to 5th year since publication), indirectly citing articles experienced a peak median harm (interquartile range) of 27.38% (-42.30% to 71.87%) in the fourth year and a minimum median harm (interquartile range) of 21.22% (-47.63% to 67.02%) in the second year. Compared to direct citations, indirect citations tend to reduce the visibility of retracted papers but are associated with more severe harm. As illustrated in Figures 3(B)-(E), the harm to the paper
indirectly citing retracted papers becomes increasingly significant as the distance of indirect citations grows. For articles connected via six degrees of separation, the maximum median harm (interquartile range) within the first fourth years reached 46.55% (-25.15% to 100%), peaking in the fourth year, while the minimum median harm (interquartile range) was 41.72% (-29.80% to 100%), observed in the second year. These findings suggest that as the citation distance increases, the link to the original retracted paper becomes harder to identify, yet the associated harm escalates. Detailed statistics on the harm experienced by indirectly citing articles are provided in Supplementary Table S3.
Notably, with greater citation distance, the proportion of duplicate articles in the citation chain also increases significantly. To more accurately assess the impact of indirect citations, we analyzed the harm to indirectly citing articles after removing duplicates. The results revealed a pattern similar to that observed before deduplication, but the harm to deduplicated articles was even greater. Detailed data for deduplicated analyses are available in Supplementary Figure S1 and Table S4.
@@
Impact of Journal Prestige on the Harm Caused by Retracted Papers.
The journal impact factor (IF) is widely used to quantify a journal’s influence. Although it does not fully capture the nuances of a journal’s impact, it remains the most widely accepted metric in academic publishing (?).
To investigate how journal impact factors influence the harm caused by retracted papers, we utilized the SciSciNet dataset (99,845,112 papers) to match all articles with their respective journals and analyzed papers published across journals with varying IFs. As shown in Figure 4(A), articles published in journals with an IF below 10 experienced significantly greater harm compared to those in journals with an IF above 10. Remarkably, for articles published in journals with an IF exceeding 20, the median harm during the first few years after publication is minimal and may even be negative.
A closer look at Figures 4(B)-(F) reveals that, while the overall harm patterns are consistent across all IF categories, notable differences appear when examining citation distances of two steps or fewer. Articles published in journals with an IF above 10 suffered significantly less harm than those published in journals with an IF below 10. In contrast, for citation distances greater than two, the impact factor of the publishing journal did not exhibit a statistically significant effect on the harm experienced by citing articles. Detailed data is available in Supplementary Table S5.
@@
Characteristics of Harm Caused by Retracted Papers Before and After Retraction.
Retraction has become an important measure to mitigate the negative impacts of problematic papers and is considered to play a critical remedial and deterrent role in curbing the publication and dissemination of such papers. To investigate the effect of retraction on the spread of academic harm, we analyzed citation patterns before and after retraction. As shown in Figure 5(A), the harm experienced by papers directly citing retracted articles is increased by citations made after the retraction. Prior to retraction, papers citing the retracted papers exhibited a median maximum harm (interquartile range) of 16.67% (-63.94% to 65.39%) within the first four years after publication. Post-retraction, the median maximum harm (interquartile range) increased to 22.28% (-59.67% to 72.98%).
In contrast, as illustrated in Figrue 5(B)-(D), the harm to papers with indirect citation distances of 2–4 is somewhat mitigated by retraction, though the effect is not pronounced. Furthermore, as depicted in Figure 5(E)-(F), retraction has a significant effect in reducing the harm to papers with indirect citation distances of 5 and 6. However, papers citing retracted articles after their retraction continue to experience considerable harm. Detailed data is available in Supplementary Table S6.
@@
Discussion
In this study, we propose an innovative measurement framework that quantifies the impact of retracted papers on the scientific community for the first time. Through a systematic analysis of the harms caused by retracted papers, we identified an ”attention escape” mechanism that enables retracted papers to persistently inflict significant negative effects on the scientific community. Specifically, our findings reveal the following characteristics of the harm caused by retracted papers: short-term effects are insignificant, but long-term effects are substantial; the impact is smaller on directly cited papers but significant on indirectly cited ones; the harm is less pronounced in high-impact journals but more significant in lower-impact journals. Our findings provide new insights into the development of more effective strategies for addressing the issues posed by retracted papers and lay the groundwork for advancing research in this field.
This study provides the first quantitative analysis of the harm caused by retracted papers and the mechanisms underlying their dissemination, uncovering specific patterns and potential drivers of their spread. While, as expected, retraction does impact the harm caused by these papers, the effects are limited. Relying solely on punitive measures, such as retraction or marking retracted citations, to curb their dissemination remains challenging. Our identification of the ”attention escape” mechanism in the harm caused by retracted papers prompts a reevaluation of strategies, suggesting a shift from penalizing stakeholders to encouraging proactive accountability through highlighting reputational and institutional risks. Our findings will drive the field to shift from merely issuing warnings to stakeholders toward awakening their intrinsic sense of responsibility. This transformation will pave the way for more comprehensive and effective approaches, strengthening research integrity and fostering a greater sense of accountability across the scientific community.

### v2

@@
Despite enormous efforts devoted to understanding the characteristics and impacts of retracted papers, little is known about the mechanisms underlying the dynamics of their harm and the dynamics of its propagation. Here, we propose a citation-based framework to quantify the harm caused by retracted papers, aiming to uncover why their harm persists and spreads so widely. We uncover an “attention escape” mechanism, wherein retracted papers postpone significant harm, more prominently affect indirectly citing papers, and inflict greater harm on citations in journals with an impact factor less than . This mechanism allows retracted papers to inflict harm outside the attention of authors and publishers, thereby evading their intervention. This study deepens understanding of the harm caused by retracted papers, emphasizes the need to activate and enhance the attention of authors and publishers, and offers new insights and a foundation for strategies to mitigate their harm and prevent its spread.
1 Introduction
@@
As of January 31, 2025, more than 60,000 papers have been publicly retracted, and this staggering number may only represent the tip of the iceberg [1, 2]. These papers have even dealt a fatal blow to entire research fields, as exemplified by the devastating cases of fraud in stem cell research [3] and Alzheimer’s research [4]. Worse still, in recent years, the number of retracted papers has remained remarkably high each year, making the challenges we face increasingly severe [5, 6, 1].
@@
Research on retracted papers has made significant progress, but most efforts have focused on their bibliometric and textual characteristics [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], whereas relatively little attention has been paid to the harm these papers cause and the mechanisms by which such harm spreads [21, 22]. Currently, researchers widely believe that the persistent harm of retracted papers stems from the difficulty authors, reviewers, and journals face in identifying and tracking their citations [23, 24, 25]. To address this issue, tools such as Feet of Clay Detector [26], Annulled Detector [27], and RetractoBot [28] have been developed to flag citations of retracted papers and alert authors to handle such references with caution. However, recent studies have suggested that these measures have been less effective than anticipated. One study found that the act of flagging citations of retracted papers provoked resistance from some authors [29]. Another study flagged 88 articles citing retracted papers and sent notification emails to the authors, but only 44 responded, and just 11 implemented corrective actions [30]. These findings indicated that current understanding of the harm caused by retracted papers remains significantly limited and biased. To mitigate the long-term impact of retracted literature, more in-depth research and fresh perspectives are urgently needed to drive meaningful progress in this field.
@@
Citations are a key and typical form through which the harm of retracted papers arises and spreads. At the same time, citations, like publications, are regarded as the “currency” of academia, forming the foundation of scholarly influence [24]. They bolster the reputation of authors and institutions, drive career advancement, secure funding opportunities, and contribute to journal rankings and impact metrics [31, 32, 33, 34]. Recognizing this dual role, we propose a citation-based quantitative framework to deepen understanding of the harm caused by retracted papers and uncover the mechanisms driving it (see “Methods” section for details). To this end, we analyzed a refined dataset of 210,309,367 publications across 23 disciplines, extracted from the Semantic Scholar dataset following data cleaning and preprocessing.(SSD, see “Methods” section) [35]. By integrating data from the Retraction Watch dataset (RWD) [7], which, after data cleaning and preprocessing, cataloged 39,449 retracted papers, we constructed an extensive citation network of retracted publications. This network enabled us to systematically trace the pathways through which harm originates and propagates, uncovering patterns that link retracted papers to their broader impact on the research community.
@@
As illustrated in Figure 1, to investigate the long-term impact of retracted papers, we focused on papers published before 2014 within the SSD, creating a subset dataset named SSD_2014 to track 10 years of citation dynamics. From the SSD_2014 dataset, we identified papers directly citing those in the RWD to form the citation distance 1 set (). Subsequently, we iteratively extracted papers citing to form the citation distance 2, 3, 4, 5, and 6 sets, denoted as , , , , and , respectively. For each of these sets, we calculated annual citation counts over 10 years and determined the expected citation counts based on contemporaneous papers published in the same journals and fields. The shortfall between observed and expected citations was defined as the harm caused by retracted papers, providing a quantitative measure of their harm and the dissemination of the harm. Details of the methodology are provided in the “Methods” section. Our research uncovered an “attention escape” mechanism by which retracted papers cause long-term, widespread, and significant harm to the scientific community. This mechanism operates through three key strategies: (1) Papers citing retracted studies experience insignificant harm during the early years when they receive the most attention, but this harm becomes increasingly significant over time. (2) Papers that indirectly cite retracted papers suffer more significant harm compared to those that directly cite them, and the severity of this harm increases with greater citation distance. (3) Papers citing retracted papers that are published in journals with an impact factor less than 10 experience more significant harm compared to those published in higher-impact, more widely recognized journals. This mechanism provides a new understanding of the harm caused by retracted papers, prompting us to shift our focus toward reactivating and enhancing the attention of authors, journals, and other stakeholders [36], thereby offering a fresh perspective for exploring new measures to mitigate the harm of retracted papers.
2 Results
@@
2.1 Temporal dynamics of harm arising from retracted papers
@@
To investigate how the harm of retracted papers evolves over time, we analyzed the annual citation impact within 10 years of publication for papers that cited these retracted works. Notably, the “10 years after publication” refers to the period starting from the second year to the eleventh year after the citing papers were published. The first year was excluded from the analysis because publication dates can vary widely throughout the year (e.g., January or December), which may result in significant biases in citation data for the first year. As shown in Figure 2(a), the harm to papers directly citing retracted papers during the first four years after publication was not significant, with the highest median (interquartile range) harm observed in the fourth year, at 18.67% (-62.92% to 68.09%). Starting in the fifth year, the harm to papers directly citing retracted papers exceeded 20%, with a median (interquartile range) of 22.84% (-57.55% to 71.98%). The lowest median (interquartile range) damage to the paper directly citing retracted papers was recorded in the second year at 10.68% (-68.71% to 62.46%). Beyond the second year, the harm to the paper directly citing retracted papers increased annually, although the rate of increase slowed over time, reaching a median (interquartile range) of 40.08% (-41.39% to 100%) by the tenth year. These findings suggest that the harm caused by retracted papers to citing papers accumulates over time, making long-term impacts unavoidable. We further analyzed the field-specific effects of retracted papers using Semantic Scholar’s discipline classifications. As shown in Figure 2(b), disciplines with more than 300 citing articles (representing 99.65% of the dataset) exhibited trends of harm to papers that directly cite retracted papers similar to the overall pattern. However, certain fields—such as computer science, physics, business, and economics (representing 2.26% of the dataset)—showed significant harm to papers that directly cite retracted papers earlier in the citing papers’ lifespans. It should be noted that a single article may belong to multiple fields, so the total percentages exceed 100%. Detailed statistics for each discipline are provided in Extended Data Table 1.
@@
Notably, in exploring the reasons behind the lack of attention to the harmful effects of problematic papers, we identified two phenomena potentially linked to the insufficient focus on the damaging mechanisms we uncovered, based on the interests most relevant to stakeholders. Two primary actors contributed to the dissemination of damage: authors (who cited retracted papers) and journals (which published papers citing retracted works). In our investigation of authors, we focused on universities as key research institutions. As shown in Table 2, among the top 30 universities in the world, we found that 28 institutions placed significant emphasis on recent research outputs (typically within the past 3 to 5 years) when considering faculty promotions [37]. In our investigation of journals, we found that most journals prioritized impact factor as a key metric of interest, with impact factor calculations typically based on papers published within the past 3 years [38]. These phenomena suggested that papers from earlier years received far less attention from stakeholders, which could have been a key reason why the potential harm of problematic papers was overlooked.
@@
2.2 Characteristics of the harm propagation arising from retracted papers
To explore how the harm from retracted papers propagates, we drew upon the “six degrees of separation” theory to study articles that were connected to retracted papers via indirect citations within six steps [39]. As shown in Figure 3(a), while the pattern of harm experienced by indirectly citing articles was similar to that of directly citing articles, the harm to indirectly citing articles was more pronounced. Specifically, within the first four years after publication, indirectly citing articles experienced a peak median harm (interquartile range) of 27.38% (-42.30% to 71.87%) in the fourth year and a minimum median harm (interquartile range) of 21.22% (-47.63% to 67.02%) in the second year. Compared to direct citations, indirect citations tended to reduce the visibility of retracted papers but were associated with more severe harm. As illustrated in Figures 3(b)-(e), the harm to the paper indirectly citing retracted papers became increasingly significant as the distance of indirect citations grew. For articles connected via six degrees of separation, within the first four years, the median harm (interquartile range) peaked at 46.55% (-25.15% to 100%) in the fourth year, while its minimum of 41.72% (-29.80% to 100%) was observed in the second year. These findings suggest that as the citation distance increases, the link to the original retracted paper becomes harder to identify, yet the associated harm escalates. Detailed statistics on the harm experienced by indirectly citing articles are provided in Extended Data Table 3.
Notably, with greater citation distance, the proportion of duplicate articles in the citation chain also increased significantly. To more accurately assess the impact of indirect citations, we analyzed the harm to indirectly citing articles after removing duplicates. The results revealed a pattern similar to that observed before deduplication, but the harm to deduplicated articles was even greater. Detailed data for deduplicated analyses are available in Extended Data Figure 1 and Table 4.
@@ 2.2 Characteristics of the harm propagation arising from retracted papers
2.3 Impact of journal prestige on the harm caused by retracted papers
The journal impact factor (IF) is widely used to quantify a journal’s influence. Although it does not fully capture the nuances of a journal’s impact, it remains the most widely accepted metric in academic publishing [40].
To investigate how journal impact factors influence the harm caused by retracted papers, we utilized the SciSciNet dataset (99,845,112 papers) to match all articles with their respective journals and analyzed papers published across journals with varying IFs. As shown in Figure 4(a), articles published in journals with an IF below 10 experienced significantly greater harm compared to those in journals with an IF above 10. Remarkably, for articles published in journals with an IF exceeding 20, the median harm during the first few years after publication was minimal and could even be negative.
A closer look at Figures 4(b)-(f) reveals that, while the overall harm patterns are consistent across all IF categories, notable differences appear when examining citation distances of two steps or fewer. Articles published in journals with an IF above 10 suffered significantly less harm than those published in journals with an IF below 10. In contrast, for citation distances greater than two, the impact factor of the publishing journal did not exhibit a statistically significant effect on the harm experienced by citing articles. Detailed data is available in Extended Data Table 5.
@@ 2.3 Impact of journal prestige on the harm caused by retracted papers
2.4 Characteristics of harm caused by retracted papers before and after retraction
Retraction has become an important measure to mitigate the negative impacts of problematic papers and is considered to play a critical remedial and deterrent role in curbing the publication and dissemination of such papers. To investigate the effect of retraction on the spread of academic harm, we analyzed citation patterns before and after retraction. As shown in Figure 5(a), the harm experienced by papers directly citing retracted articles increased by citations made after the retraction. Prior to retraction, papers citing the retracted papers exhibited a median maximum harm (interquartile range) of 16.67% (-63.94% to 65.39%) within the first four years after publication. After Retraction, the median maximum harm (interquartile range) increased to 22.28% (-59.67% to 72.98%).
In contrast, as illustrated in Figrue 5(b)-(d), the harm to papers with indirect citation distances of 2–4 was somewhat mitigated by retraction, though the effect was not pronounced. Furthermore, as depicted in Figure 5(e)-(f), retraction had a significant effect in reducing the harm to papers with indirect citation distances of 5 and 6. However, papers citing retracted articles after their retraction continued to experience considerable harm. Detailed data is available in Extended Data Table 6.
@@ 2.4 Characteristics of harm caused by retracted papers before and after retraction
3 Discussion
In this study, we propose an innovative measurement framework that quantifies the impact of retracted papers on the scientific community for the first time. Through a systematic analysis of the harms caused by retracted papers, we identify an “attention escape” mechanism that enables retracted papers to persistently inflict significant negative effects on the scientific community. Specifically, our findings reveal the following characteristics of the harm caused by retracted papers: short-term effects are insignificant, but long-term effects are substantial; the impact is smaller on directly cited papers but significant on indirectly cited ones; the harm is less pronounced in high-impact journals but more significant in lower-impact journals. Our findings provide new insights into the development of more effective strategies for addressing the issues posed by retracted papers and lay the groundwork for advancing research in this field.
This study provides the first quantitative analysis of the harm caused by retracted papers and the mechanisms underlying their dissemination, uncovering specific patterns and potential drivers of their spread. While, as expected, retraction does impact the harm caused by these papers, the effects are limited. Relying solely on punitive measures, such as retraction or marking retracted citations, to curb their dissemination remains challenging. Our identification of the “attention escape” mechanism in the harm caused by retracted papers prompts a reevaluation of strategies, suggesting a shift from penalizing stakeholders to encouraging proactive accountability through highlighting reputational and institutional risks. Our findings will drive the field to shift from merely issuing warnings to stakeholders toward awakening their intrinsic sense of responsibility. This transformation will pave the way for more comprehensive and effective approaches, strengthening research integrity and fostering a greater sense of accountability across the scientific community.